OpenStudy (anonymous):

What is the quotient: (64y3 + 20y2 – 28y) ÷ 4y ? 16y2 – 5y + 7 16y3 + 5y2 – 7y 16y2 + 5y – 7 16y2 – 5y – 7

OpenStudy (lgbasallote):

try separating the numerator: \[\implies \frac{64y^3}{4y} + \frac{20y^2}{4y} - \frac{28y}{4y}\] does that help?

OpenStudy (anonymous):

how do i get 16?

OpenStudy (anonymous):

oh, i divide it

OpenStudy (lgbasallote):

yes. very perceptive

OpenStudy (anonymous):

would it be 16y 2 or 3?

OpenStudy (anonymous):

B?

OpenStudy (lgbasallote):

what is y^3 divided by y? y^2 or y^3?

OpenStudy (anonymous):

3rd?

OpenStudy (lgbasallote):

yes